# Realising VCD for untwisted automorphism groups of RAAGs

### v1

@@ 4.4. The retraction process
In this and the next subsection we prove Theorem D (i), restated below as Theorem 4.16.
@@
We thank Karen Vogtmann for pointing out this clarification.
Fix $\Gamma$ spiky. In particular, both Lemma 4.10 and Lemma 4.13 apply.
@@
$C$$F$push into $C$along $F$
Figure 9. Pushing along $F$ into $C$, turning $C$ from a 3-dimensional cube to a 2-dimensional cube complex looking like an open box or shell. The red arrows show how we visualise pushing in $F$.
In fact, we will do slightly more radical deformation retractions than this. We will prove that the structure of $K_{\Gamma}$ is such that if $F$ and $F^{\prime}$ are both free faces of $C$, then so is $F\cap F^{\prime}$. We will identify many free faces of $C$, and prove that we can push along their (non-empty) common intersection into $C$. By push along a subface $F$ into $C$, we mean a deformation retraction throughout which every cube not containing $F$ is fixed, while the interior of every cube containing $F$ is pushed onto this fixed subcomplex, as demonstrated below in Figure 10.
@@
[…]
We proceed by finding free faces of $K_{\Gamma}$, retracting these as described (and doing so equivariantly with respect to the $U(A_{\Gamma})$-action). Then we look for more free faces in the resultant complex; when we can no longer find free faces, we stop. If we manage to retract all cubes of dimension greater than some $k$, then the complex obtained at the end of this process will have definition at most $k$.
We identify free faces by examining the structure of compatible sets of $\Gamma$-partitions with respect to the hugged partitions they contain. Intuitively, if a set $\Pi$ of partitions contains hugged partitions (or if non-principal partitions not in $\Pi$ would be hugged in $\Pi$ upon being added to $\Pi$), then $\Pi$ contains some redundant information. Partitions which are hugged in $\Pi$ contribute roughly the same structure to $K_{\Gamma}$ as the principal partitions which hug them—and this turns out to result in $K_{\Gamma}$ having many free faces. Our procedure stops when we have no more hugged partitions, so we have a characterisation of the final complex in terms of hugged partitions (which we state in Proposition 4.19).
@@
The outermost (i.e. the slowest-moving) variable in this induction is $b$, followed by $t$ and then by $p$. We will therefore start with $b=0$, that is, by considering cubes whose bottom vertex is a marked Salvetti. Working at a fixed value of $b$, we begin with $t=0$, that is, with cubes whose top vertex corresponds to a blowup by a set of partitions of maximal possible size, $M(V)$. Finally, for each fixed pair $(b,t)$, we start with $p=0$—that is, among cubes $c(\Pi_{1},\Pi_{2})$ with $|\Pi_{1}|=b$ and $M(V)-|\Pi_{2}|=t$, we begin by considering those $\Pi_{2}$ containing the maximal possible number of principal partitions.
@@
We examine which partitions could be added to the set $\Pi_{2}\setminus H$. We cannot add any principal partitions to it, since $\Pi_{2}\setminus H$ already contains $M(L)$ principal partitions. On the other hand, by Lemma 4.10, any non-principal partition compatible with $\Pi_{2}\setminus H$ would necessarily be compatible with all of $H$ too, so since $\Pi_{2}$ was maximally-sized, any such is an element of $H$. In other words, the only partitions we may add to $\Pi_{2}\setminus H$ are elements of $H$, so the only cubes containing $c(\emptyset,\Pi_{2}\setminus H)$ are contained in $c(\emptyset,\Pi_{2})$. Since $c(\emptyset,\Pi_{2}\setminus H;\alpha)$ has a vertex at a Salvetti complex with marking $\alpha$, any other cube containing it as a subface necessarily also has marking $\alpha$. We can now perform a deformation retraction by pushing into the cube $c(\emptyset,\Pi_{2})$ along its free subface $c(\emptyset,\Pi_{2}\setminus H)$. Doing this simultaneously at all marked Salvetti complexes is a $U(A_{\Gamma})$-equivariant retraction, since $U(A_{\Gamma})$ acts only by changing markings. We do the same retractions for all cubes $c(\emptyset,\Pi_{2})$ with $\Pi_{2}$ having parameters $t=0,\;p=0$, which completes the $p=0$ base case.
Case Ib: $b=0$$t=0$$p=k>0$.
@@
Our inductive hypothesis is that all cases $t=0,\;p<k$ have been handled in the above way: all hugged partitions were removed, and then the retraction was effected by pushing in along this subface. Now fix $\Pi_{2}$ with $t=0$ and $p=k$; let $H$ be its set of hugged partitions and consider $\Pi\setminus H$. Once again, we wish to prove that $c(\emptyset,\Pi_{2}\setminus H)$ is a free face of $c(\emptyset,\Pi_{2})$.
@@
If $\mathcal{P}$ is a principal partition which is compatible with all of $\Pi_{2}\setminus H$, then by Corollary 4.14, and since $\Pi_{2}$ is maximally-sized, there is exactly one $\mathcal{Q}\in H$ with which $\mathcal{P}$ is not compatible. Thus we may form the cube $c(\emptyset,(\Pi_{2}\setminus\mathcal{Q})\cup\mathcal{P})$. This falls under the case $t=0,\;p=k-1$, so by the inductive hypothesis, has already been retracted. Moreover, it was retracted via the subface obtained by removing all of its hugged partitions—which in particular involves removing all of $H\setminus\mathcal{Q}$. Hence, the cube $c(\emptyset,(\Pi_{2}\setminus H)\cup\mathcal{P})$ has already been removed.
@@
Instead let $\mathcal{Q}^{\prime}$ be a non-principal partition which is compatible with all of $\Pi_{2}\setminus H$. Then by Lemma 4.10 it is compatible with all of $H$; since $t=0$ (that is, $\Pi_{2}$ is maximally-sized), we must have $\mathcal{Q}^{\prime}\in H$. Therefore, for each marking $\alpha$, the only cubes containing $c(\emptyset,\Pi_{2}\setminus H;\alpha)$ are themselves contained in $c(\emptyset,\Pi_{2};\alpha)$. Hence, for each marking $\alpha$, we may retract $c(\emptyset,\Pi_{2};\alpha)$ by pushing in along $c(\emptyset,\Pi_{2}\setminus H;\alpha)$; this is once again $U(A_{\Gamma})$-equivariant. Do this simultaneously for all $\Pi_{2}$ with $t=0,\;p=k$; doing this for each possible $p=k>0~{}$ completes the $t=0$ step.
Case Ic: $b=0$$t=l>0$$p=k\geq 0$.
@@
Let $\mathcal{P}$ be a principal partition compatible with all of $\Pi_{2}\setminus H$. Suppose first that $\mathcal{P}$ is also compatible with all of $H$. Then the set $\Pi_{2}\cup\mathcal{P}$ of pairwise compatible partitions has parameters $t=l-1,\;p=k$, so by the inductive hypothesis, has already been retracted. This retraction happened by removing all the hugged partitions in $\Pi_{2}\cup\mathcal{P}$, which in particular includes all of $H$. Hence the cube $c(\emptyset,(\Pi_{2}\setminus H)\cup\mathcal{P})$ is no longer present in our complex.
Instead suppose that $\mathcal{P}$ is not compatible with all of $H$, so by Lemma 4.13 there is exactly one $\mathcal{Q}\in H$ with which $\mathcal{P}$ is not compatible. We can therefore replace $\mathcal{Q}$ with $\mathcal{P}$, and obtain the compatible set $(\Pi_{2}\setminus\mathcal{Q})\cup\mathcal{P}$. This has parameters $t=l,\;p=k-1$, and so the cube $c(\emptyset,(\Pi_{2}\setminus H)\cup\mathcal{P})$ has also already been dealt with previously in the process. The set of hugged partitions in this case contains $H\setminus\mathcal{Q}$, so the subface used to retract this cube was contained in $c(\emptyset,(\Pi_{2}\setminus H)\cup\mathcal{P})$, so we have already removed $c(\emptyset,(\Pi_{2}\setminus H)\cup\mathcal{P})$.
@@
We are left in the case $\mathcal{Q}^{\prime}\in H$—now the cube $c(\emptyset,(\Pi_{2}\setminus H)\cup\mathcal{Q}^{\prime})$ is contained in $c(\emptyset,\Pi_{2})$ and will get retracted when we push in along $c(\emptyset,\Pi_{2}\setminus H)$.
Hence, at this stage of our retraction process, the only cubes which contain $c(\emptyset,\Pi_{2}\setminus H)$ are themselves contained in $c(\emptyset,\Pi_{2})$ so we may push into $c(\emptyset,\Pi_{2})$ along $c(\emptyset,\Pi_{2}\setminus H)$. Once again, we do this simultaneously for every marking; this is again $U(A_{\Gamma})$-equivariant. We may do this simultaneously for all cubes $c(\emptyset,\Pi_{2})$ where $\Pi_{2}$ has parameters $t=l,\;p=k$, which completes the induction step.
@@
    For all markings $\beta$, we have also retracted all cubes of the form $c(\Pi_{1}^{\prime},\Pi_{2}^{\prime};\beta)$ where $|\Pi_{1}^{\prime}|=b$ and where $c(\emptyset,\Pi_{2}^{\prime})$ would have been retracted before $c(\emptyset,\Pi_{2})$ in the $b=0$ step (that is, $t(\Pi_{2}^{\prime})<t(\Pi_{2})$, or $t(\Pi_{2}^{\prime})=t(\Pi_{2})$ and $p(\Pi_{2}^{\prime})<p(\Pi_{2})$).
    Moreover, we assume that in each case the retraction has been done via the subface $c(\Pi_{1}^{\prime},\Pi_{2}^{\prime}\setminus H^{\prime};\beta)$, where $H^{\prime}$ is the set of non-principal partitions in $\Pi_{2}^{\prime}\setminus\Pi_{1}^{\prime}$ which are hugged in $\Pi_{2}^{\prime}$.
@@
We could remove some partition $\mathcal{R}$ from $\Pi_{1}$, yielding a subface $c(\Pi_{1}\setminus\mathcal{R},\Pi_{2}\setminus H;\alpha)$. Now, if $\mathcal{R}$ was hugged in $\Pi_{2}$, then the cube $c(\Pi_{1}\setminus\mathcal{R},\Pi_{2}\setminus(H\cup\mathcal{R});\alpha)$ was used to retract $c(\Pi_{1}\setminus\mathcal{R},\Pi_{2};\alpha)$ in the $b=m-1$ case (using part (I) of the inductive hypothesis). In particular, this retraction will also have removed $c(\Pi_{1},\Pi_{2};\alpha)$, so we may ignore this case. If instead $\mathcal{R}$ was not hugged in $\Pi_{2}$, then $c(\Pi_{1}\setminus\mathcal{R},\Pi_{2};\alpha)$ was retracted by pushing in along $c(\Pi_{1}\setminus\mathcal{R},\Pi_{2}\setminus H;\alpha)$. In either case, we need not consider the cube $c(\Pi_{1}\setminus\mathcal{R},\Pi_{2}\setminus H;\alpha)$.
Instead, we could add a partition $\mathcal{R}$ to $\Pi_{2}$, to get the cube $c(\Pi_{1},(\Pi_{2}\setminus H)\cup\mathcal{R};\alpha)$. We can ignore the case $\mathcal{R}\in H$, since pushing in to $c(\Pi_{1},\Pi_{2};\alpha)$ along $c(\Pi_{1},\Pi_{2}\setminus H;\alpha)$ will remove the face $c(\Pi_{1},(\Pi_{2}\setminus H)\cup\mathcal{R};\alpha)$ anyway. So now by (II), the cube $c(\Pi_{1},\Pi_{2}\cup\mathcal{R};\alpha)$ has already been retracted (since it corresponds to the case $t=k-1$). This retraction used the subface obtained by removing all the non-principal partitions in $(\Pi_{2}\cup\mathcal{R})\setminus\Pi_{1}$ which are hugged in $\Pi_{2}\cup\mathcal{R}$; in particular, this includes removing all of $H$. Hence, the cube $c(\Pi_{1},(\Pi_{2}\setminus H)\cup\mathcal{R};\alpha)$ has already been removed in a previous retraction. (Once again, if $\mathcal{R}$ is (non-principal and) hugged in $\Pi_{2}$, then we may move on, as the cube $c(\Pi_{1},\Pi_{2};\alpha)$ was also removed in this retraction.)
Therefore, the only cubes with marking $\alpha$ which contain $c(\Pi_{1},\Pi_{2}\setminus H;\alpha)$ are themselves contained in $c(\Pi_{1},\Pi_{2};\alpha)$. We now turn our attention to cubes with other markings. We have the following general lemma.
We say that a compatible set of $\Gamma$-partitions $\Pi$ is inextendible if no $\Gamma$-partitions can be added to $\Pi$. That is, there are no $\Gamma$-partitions not in $\Pi$ which are compatible with all elements of $\Pi$.
@@
Let $\Gamma$ be any simplicial graph. Let $\Pi$ be any inextendible set of $\Gamma$-partitions. Let $\alpha$$\beta$ be (distinct, non-equivalent) markings of the $\Gamma$-complex $\mathbb{S}_{\Gamma}^{\Pi}$, and suppose that the cubes $\mathcal{C}_{\alpha}:=c(\emptyset,\Pi;\alpha)$ and $\mathcal{C}_{\beta}:=c(\emptyset,\Pi;\beta)$ meet along some common face. Then this face is contained in a face common to both cubes which contains the vertex corresponding to the maximal blowup, $\mathbb{S}_{\Gamma}^{\Pi}$.

### v2

@@ 4.4. The retraction process
In this and the next subsection we prove Theorem D (i), which we restate as follows.
Theorem 4.16.
@@
Fix $\Gamma$ spiky. In particular, both our key lemmas (4.11 & 4.14) apply.
@@
CCFFpush along FFinto $C$
Figure 9. Pushing along $F$ into $C$, turning $C$ from a 3-dimensional cube to a 2-dimensional cube complex looking like an open box or shell. The (red) dotted arrows show how we visualise pushing in $F$.
In fact, we will do slightly more radical deformation retractions than this. We will prove that the structure of $K_{\Gamma}$ is such that if $F$ and $F^{\prime}$ are both free faces of $C$, then so is $F\cap F^{\prime}$ (observe that this is not true for arbitrary cube complexes: two cubes could ‘hinge’ along a common codimension-2 subface, for instance). We will identify many free faces of $C$, and prove that we can push along their (non-empty) common intersection into $C$. By push along a subface into $C$, we mean a deformation retraction in which every cube not containing this subface is fixed, while the interior of every cube containing the subface is pushed onto this fixed subcomplex, as demonstrated in Figure 10.
@@
Figure 10. Here we are pushing along the intersection $F\cap F^{\prime}$ into $C$. The interiors of the faces $^{\prime}FFF\cap F^{\prime}$, as well as the interior of $C$, are all pushed onto the 2-dimensional subcomplex shown on the right.
We proceed by finding free faces of $K_{\Gamma}$, retracting these as described (and doing so equivariantly). Then we look for more free faces in the resultant complex; when we can no longer find free faces, we stop. If we have managed to retract all cubes of dimension greater than some $k$, then the complex we have obtained has dimension at most $k$.
We identify free faces by examining the structure of compatible sets of $\Gamma$-partitions with respect to the hugged partitions they contain. Intuitively, if a set $\Pi$ of partitions contains hugged partitions (or if non-principal partitions not in $\Pi$ would be hugged in $\Pi$ upon being added to $\Pi$), then $\Pi$ contains some redundant information. Partitions which are hugged in $\Pi$ contribute roughly the same structure to $K_{\Gamma}$ as the principal partitions which hug them – and this turns out to result in $K_{\Gamma}$ having many free faces. Our procedure stops when we have no more hugged partitions, so we have a characterisation of the final complex in terms of hugged partitions (which we state in Proposition 4.20).
@@
The outermost (i.e. the slowest-moving) variable in this induction is $b$, followed by $t$ and then by $p$. We will therefore start with $b=0$, that is, by considering cubes whose bottom vertex is a marked Salvetti. Working at a fixed value of $b$, we begin with $t=0$, that is, with cubes whose top vertex corresponds to a blowup by a set of partitions of maximal possible size, $M(V)$. Finally, for each fixed pair $(b,t)$, we start with $p=0$ – that is, among cubes $c(\Pi_{1},\Pi_{2})$ with $|\Pi_{1}|=b$ and $M(V)-|\Pi_{2}|=t$, we begin by considering those $\Pi_{2}$ containing the maximal possible number of principal partitions.
@@
We examine which partitions could be added to the set $\Pi_{2}\setminus H$. We cannot add any principal partitions to it, since $\Pi_{2}\setminus H$ already contains $M(L)$ principal partitions. On the other hand, by Lemma 4.11, any non-principal partition compatible with $\Pi_{2}\setminus H$ would necessarily be compatible with all of $H$ too, so since $\Pi_{2}$ was maximally-sized, any such is an element of $H$. In other words, the only partitions we may add to $\Pi_{2}\setminus H$ are elements of $H$, so the only cubes containing $c(\emptyset,\Pi_{2}\setminus H)$ are contained in $c(\emptyset,\Pi_{2})$. Since $c(\emptyset,\Pi_{2}\setminus H;\alpha)$ has a vertex at a Salvetti complex with marking $\alpha$, any other cube containing it as a subface necessarily also has marking $\alpha$. We can now perform a deformation retraction by pushing into the cube $c(\emptyset,\Pi_{2})$ along its free subface $c(\emptyset,\Pi_{2}\setminus H)$. Doing this simultaneously at all marked Salvetti complexes is a $U(A_{\Gamma})$-equivariant retraction, since $U(A_{\Gamma})$ acts only by changing markings. We do the same retractions for all cubes $c(\emptyset,\Pi_{2})$ with $\Pi_{2}$ having parameters t=0,p=0t=0,\;p=0, which completes the $p=0$ base case.
Case Ib: b=0b=0t=0t=0p=k>0p=k>0.
@@
Our inductive hypothesis is that all cases t=0,p<kt=0,\;p<k have been handled in the above way: all hugged partitions were removed, and then the retraction was effected by pushing in along this subface. Now fix $\Pi_{2}$ with $t=0$ and $p=k$; let $H$ be its set of hugged partitions and consider $\Pi\setminus H$. Once again, we wish to prove that $c(\emptyset,\Pi_{2}\setminus H)$ is a free face of $c(\emptyset,\Pi_{2})$.
@@
If $\mathcal{P}$ is a principal partition which is compatible with all of $\Pi_{2}\setminus H$, then by Corollary 4.15, and since $\Pi_{2}$ is maximally-sized, there is exactly one $\mathcal{Q}\in H$ with which $\mathcal{P}$ is not compatible. Thus we may form the cube $c(\emptyset,(\Pi_{2}\setminus\mathcal{Q})\cup\mathcal{P})$. This falls under the case t=0,p=k1t=0,\;p=k-1, so by the inductive hypothesis, has already been retracted. Moreover, it was retracted via the subface obtained by removing all of its hugged partitions – which in particular involves removing all of $H\setminus\mathcal{Q}$. Hence, the cube $c(\emptyset,(\Pi_{2}\setminus H)\cup\mathcal{P})$ has already been removed.
@@
Instead let $\mathcal{Q}^{\prime}$ be a non-principal partition which is compatible with all of $\Pi_{2}\setminus H$. Then by Lemma 4.11 it is compatible with all of $H$; since $t=0$ (that is, $\Pi_{2}$ is maximally-sized), we must have $\mathcal{Q}^{\prime}\in H$. Therefore, for each marking $\alpha$, the only cubes containing $c(\emptyset,\Pi_{2}\setminus H;\alpha)$ are themselves contained in $c(\emptyset,\Pi_{2};\alpha)$. Hence, for each marking $\alpha$, we may retract $c(\emptyset,\Pi_{2};\alpha)$ by pushing in along $c(\emptyset,\Pi_{2}\setminus H;\alpha)$; this is once again $U(A_{\Gamma})$-equivariant. Do this simultaneously for all $\Pi_{2}$ with t=0,p=kt=0,\;p=k; doing this for each possible p=k>0p=k>0~ completes the $t=0$ step.
[…]
@@
Let $\mathcal{P}$ be a principal partition compatible with all of $\Pi_{2}\setminus H$. Suppose first that $\mathcal{P}$ is also compatible with all of $H$. Then the set $\Pi_{2}\cup\mathcal{P}$ of pairwise compatible partitions has parameters t=l1,p=kt=l-1,\;p=k, so by the inductive hypothesis, has already been retracted. This retraction happened by removing all the hugged partitions in $\Pi_{2}\cup\mathcal{P}$, which in particular includes all of $H$. Hence the cube $c(\emptyset,(\Pi_{2}\setminus H)\cup\mathcal{P})$ is no longer present in our complex.
Instead suppose that $\mathcal{P}$ is not compatible with all of $H$, so by Lemma 4.14 there is exactly one $\mathcal{Q}\in H$ with which $\mathcal{P}$ is not compatible. We can therefore replace $\mathcal{Q}$ with $\mathcal{P}$, and obtain the compatible set $(\Pi_{2}\setminus\mathcal{Q})\cup\mathcal{P}$. This has parameters t=l,p=k1t=l,\;p=k-1, and so the cube $c(\emptyset,(\Pi_{2}\setminus H)\cup\mathcal{P})$ has also already been dealt with previously in the process. The set of hugged partitions in this case contains $H\setminus\mathcal{Q}$, so the subface used to retract this cube was contained in $c(\emptyset,(\Pi_{2}\setminus H)\cup\mathcal{P})$, so we have already removed $c(\emptyset,(\Pi_{2}\setminus H)\cup\mathcal{P})$.
@@
We are left in the case $\mathcal{Q}^{\prime}\in H$ – now the cube $c(\emptyset,(\Pi_{2}\setminus H)\cup\mathcal{Q}^{\prime})$ is contained in $c(\emptyset,\Pi_{2})$ and will get retracted when we push in along $c(\emptyset,\Pi_{2}\setminus H)$.
Hence, at this stage of our retraction process, the only cubes which contain $c(\emptyset,\Pi_{2}\setminus H)$ are themselves contained in $c(\emptyset,\Pi_{2})$ – so we may push into $c(\emptyset,\Pi_{2})$ along $c(\emptyset,\Pi_{2}\setminus H)$. Once again, we do this simultaneously for every marking; this is again $U(A_{\Gamma})$-equivariant. We may do this simultaneously for all cubes $c(\emptyset,\Pi_{2})$ where $\Pi_{2}$ has parameters t=l,p=kt=l,\;p=k, which completes the induction step.
@@
    For all markings $\beta$, we have also retracted all cubes of the form $c(\Pi_{1}^{\prime},\Pi_{2}^{\prime};\beta)$ where we have $|\Pi_{1}^{\prime}|=b$ and where $c(\emptyset,\Pi_{2}^{\prime})$ would have been retracted before $c(\emptyset,\Pi_{2})$ in the $b=0$ step (that is, $t(\Pi_{2}^{\prime})<t(\Pi_{2})$, or $t(\Pi_{2}^{\prime})=t(\Pi_{2})$ and $p(\Pi_{2}^{\prime})<p(\Pi_{2})$).
    Also assume that in each case, retraction has been done via the subface $c(\Pi_{1}^{\prime},\Pi_{2}^{\prime}\setminus H^{\prime};\beta)$, where $H^{\prime}$ is the set of non-principal partitions in $\Pi_{2}^{\prime}\setminus\Pi_{1}^{\prime}$ which are hugged in $\Pi_{2}^{\prime}$.
@@
We could remove some partition $\mathcal{R}$ from $\Pi_{1}$, yielding a subface $c(\Pi_{1}\setminus\mathcal{R},\Pi_{2}\setminus H;\alpha)$. Now, if $\mathcal{R}$ was hugged in $\Pi_{2}$, then the cube $c(\Pi_{1}\setminus\mathcal{R},\Pi_{2}\setminus(H\cup\mathcal{R});\alpha)$ was used to retract $c(\Pi_{1}\setminus\mathcal{R},\Pi_{2};\alpha)$ in the b=m1b=m-1 case (using part (I) of the inductive hypothesis). In particular, this retraction will also have removed $c(\Pi_{1},\Pi_{2};\alpha)$, so we may ignore this case. If instead $\mathcal{R}$ was not hugged in $\Pi_{2}$, then $c(\Pi_{1}\setminus\mathcal{R},\Pi_{2};\alpha)$ was retracted by pushing in along $c(\Pi_{1}\setminus\mathcal{R},\Pi_{2}\setminus H;\alpha)$. In either case, we need not consider the cube $c(\Pi_{1}\setminus\mathcal{R},\Pi_{2}\setminus H;\alpha)$.
Instead, we could add a partition $\mathcal{R}$ to $\Pi_{2}$, to get the cube $c(\Pi_{1},(\Pi_{2}\setminus H)\cup\mathcal{R};\alpha)$. We can ignore the case $\mathcal{R}\in H$, since pushing in to $c(\Pi_{1},\Pi_{2};\alpha)$ along $c(\Pi_{1},\Pi_{2}\setminus H;\alpha)$ would remove the face $c(\Pi_{1},(\Pi_{2}\setminus H)\cup\mathcal{R};\alpha)$ anyway. So now by (II), the cube $c(\Pi_{1},\Pi_{2}\cup\mathcal{R};\alpha)$ has already been retracted (since it corresponds to the case t=k1t=k-1). This retraction used the subface obtained by removing all the non-principal partitions in $(\Pi_{2}\cup\mathcal{R})\setminus\Pi_{1}$ which are hugged in $\Pi_{2}\cup\mathcal{R}$; in particular, this includes removing all of $H$. Hence, the cube $c(\Pi_{1},(\Pi_{2}\setminus H)\cup\mathcal{R};\alpha)$ has already been removed in a previous retraction. (Once again, if $\mathcal{R}$ is (non-principal and) hugged in $\Pi_{2}$, then we may move on, as the cube $c(\Pi_{1},\Pi_{2};\alpha)$ was also removed in this retraction.)
Therefore, the only cubes with marking $\alpha$ which contain $c(\Pi_{1},\Pi_{2}\setminus H;\alpha)$ are themselves contained in $c(\Pi_{1},\Pi_{2};\alpha)$. We now turn our attention to cubes with other markings. We have the following general lemma, which we need some terminology to state: we say that a compatible set of $\Gamma$-partitions $\Pi$ is inextendible if no $\Gamma$-partitions can be added to $\Pi$. That is, there are no $\Gamma$-partitions not in $\Pi$ which are compatible with all elements of $\Pi$.
@@
Let $\Gamma$ be a simplicial graph and $\Pi$ be an inextendible set of $\Gamma$-partitions. Let $\alpha$$\beta$ be (distinct, non-equivalent) markings of the $\Gamma$-complex $\mathbb{S}_{\Gamma}^{\Pi}$. Suppose that the cubes $\mathcal{C}_{\alpha}:=c(\emptyset,\Pi;\alpha)$ and $\mathcal{C}_{\beta}:=c(\emptyset,\Pi;\beta)$ meet along some common face. Then this face is contained in a face common to both cubes which contains the vertex corresponding to the maximal blowup, $\mathbb{S}_{\Gamma}^{\Pi}$.